# MultiNash-PF: A Particle Filtering Approach for Computing Multiple Local Generalized Nash Equilibria in Trajectory Games

### v1

@@ Abstract
Modern-world robotics involves complex environments where multiple autonomous agents must interact with each other and other humans. This necessitates advanced interactive multi-agent motion planning techniques. Generalized Nash equilibrium(GNE), a solution concept in constrained game theory, provides a mathematical model to predict the outcome of interactive motion planning, where each agent needs to account for other agents in the environment. However, in practice, multiple local GNEs may exist. Finding a single GNE itself is complex as it requires solving coupled constrained optimal control problems. Furthermore, finding all such local GNEs requires exploring the solution space of GNEs, which is a challenging task. This work proposes the MultiNash-PF framework to efficiently compute multiple local GNEs in constrained trajectory games. Potential games are a class of games for which a local GNE of a trajectory game can be found by solving a single constrained optimal control problem. We propose MultiNash-PF that integrates the potential game approach with implicit particle filtering, a sample-efficient method for non-convex trajectory optimization. We first formulate the underlying game as a constrained potential game and then utilize the implicit particle filtering to identify the coarse estimates of multiple local minimizers of the game’s potential function. MultiNash-PF then refines these estimates with optimization solvers, obtaining different local GNEs. We show through numerical simulations that MultiNash-PF reduces computation time by up to 50% compared to a baseline approach.
@@ I Introduction
Modern-world robotics applications are inherently multi-agent in nature. For instance, autonomous cars on highways need to account for other vehicles while planning their motion. Many robotics applications require motion planning for selfish agents sharing a common operational space, such as autonomous cars and delivery robots. In such settings, autonomous agents should account for other agents in the shared environment while achieving their individual goals. Autonomous agents need to reason about the impact of their actions on other agents. Game theory provides a formal way to deal with such interactive scenarios [1]. Furthermore, recent works have shown that game-theoretic trajectory planning makes robots more socially acceptable to humans [2].
@@
To model the outcome of such interactive scenarios, Nash equilibrium is a key concept in game theory. When multiple agents- such as robots and humans- interact, a Nash equilibrium corresponds to a joint action such that no agent has incentives to deviate from its equilibrium actions unilaterally. This allows us to account for agents’ individual preferences when interacting in a common environment. We formulate a multi-agent interactive trajectory planning problem as a trajectory game. In robotics, seeking Nash equilibria of the underlying trajectory games has been proven to be effective [3, 4]. Furthermore, in real-world applications, agents may have to satisfy some safety constraints, such as avoiding collisions with obstacles and other agents in the environment. In such constrained multi-agent interactive settings, the interaction outcomes are captured through generalized Nash equilibrium(GNE), which are constrained Nash equilibria. Several recent works have shown great success in utilizing GNE for multi-agent motion planning in interactive domains [5, 6].
However, in practice, multiple local GNEs may exist for a trajectory game. For example, in the two-player trajectory game illustrated in Fig. 1, the agents need to move towards their goals while avoiding collisions. Here, two local GNEs exist: clockwise trajectories and counterclockwise trajectories. If the two agents were to act independently, they could reach different GNEs, which would result in a collision. Therefore, agents need a coordination mechanism to avoid such scenarios. However, identifying all the possible local GNEs is required to enable effective coordination. Therefore, computing multiple local GNEs is critical to aligning players’ preferences over different equilibria [7] or helping coordinate players’ actions via recommendation [8]. Furthermore, different local GNEs correspond to different modes of interactions in multi-agent interactive settings. Therefore, it is essential to efficiently recover all the possible multi-modal outcomes of multi-agent interactions.
@@
Computing different local GNEs requires an efficient way to solve games. In practice, computing even a single GNE is often challenging [9] because solving for GNE amounts to solving a set of coupled constrained optimal control problems. Typically, such a solution requires solving first-order KKT conditions of the underlying problems. Therefore, solving the GNE often reduces to addressing a set of complementarity conditions, a process known to be computationally demanding, especially for large-scale applications [10]. Furthermore, computing all the different local GNEs requires initializing the game solvers with many different initial guesses of the equilibrium trajectories from prior knowledge of the specific scenario. Further, as the number of participating agents increases, the dimension of the nonconvex optimization grows to interactable levels. This further increases the required number of initial guesses for the game solvers to find all local GNEs. To our best knowledge, there is no principled approach to compute multiple local Nash equilibria in constrained trajectory games other than exhaustive random initializations of the game solvers.
@@
We propose a novel framework named MultiNash-PF to efficiently compute multiple local GNEs in constrained trajectory games. The approach combines
  1. 1.
    constrained potential games [11], a class of games for which we can compute local Nash equilibria by optimizing a single constrained optimal control problem, and,
  2. 2.
    constraint aware implicit particle filtering, which is a sample-efficient particle filtering method recently used for nonconvex trajectory optimization [12].
Using constrained potential games, it is possible to compute a local GNE by solving a constrained optimal control problem using off-the-shelf optimization solvers. These solvers outperform state-of-the-art constrained dynamic game solvers in computation speed during [13]. After formulating the constrained trajectory game as a constrained potential game, our approach employs the constraint-aware implicit particle filtering method to identify multiple local minima of the potential game. This method eliminates the requirements of random initialization of the optimization solver by efficiently obtaining coarse estimates of different local GNEs. Next, we use the identified minima as initialization for the optimization solver. The solver then computes all the different local GNEs of the original game. The framework of MultiNash-PF is depicted in Fig. 1. Leveraging the sample efficiency of the implicit particle filtering method, MultiNash-PF can explore and identify the multiple local GNEs faster than the random initialization of optimization solvers.
@@ II Related Works
Game Theoretic Planning. One of the earliest works laying the foundation of game theoretic motion planning was [14]. Game theoretic motion planners that seek Stackelberg equilibria were proposed in [15, 3]. Later, methods that seek Nash equilibria were developed. These methods differ mainly in the way they compute Nash equilibria. Iterative best response methods for computing Nash equilibria were developed in [16, 17]. Iterative linear quadratic approximations to compute Nash equilibria were introduced in [18]. ALGames was developed in [19] which is an augmented lagrangian-based solver for constrained dynamic games. A sequential linear-quadratic technique game-based technique to compute GNE was proposed in [20]. However, these methods are computationally expensive for real-time implementations as the number of agents increases. Potential games-based methods were developed later in [11, 21, 13, 22] to reduce the complexity of computing equilibria. However, the shortcomings of all these methods lie in the fact that they find only one Nash equilibrium of the underlying game and ignore the multi-modal nature of the problem that is crucial for coordination among agents. In [7], the authors recognize the existence of multiple equilibria and focus on inferring the equilibrium of other agents in the environment. However, the methodology relies on computing the different approximate local Nash by rollouts of randomly selected initial strategies, which we show in this work to be computationally intractable with the risk of not identifying all possible equilibria.
@@
Multiple Local Solutions to Constrained Optimal Control. In general, optimal control problems are often formulated as nonconvex constrained optimization problems, which are solved through the conventional gradient-based approaches such as inter-point methods [23] and sequential quadratic programming [24]. However, given an initialization, they can only find a single local optimal solution. However, in the present context of identifying all potential local GNEs, the gradient-based method would require numerous initializations to find all the local solutions. Departing from gradient-based methods, a Bayesian inference approach has been proposed to solve model predictive control problem [25, 26, 27]. The approach formulates the optimal control problem from a Bayesian inference perspective and employs particle filtering as its estimation method. At its core, particle filtering utilizes sequential Monte Carlo sampling to perform state estimation for highly nonlinear systems [28, 29]. The efficiency and accuracy of the estimation largely depend on the quality of the sampled particles. Recently, in [30], an efficient particle filtering method was proposed based on the principle of implicit sampling. This sampling approach relies on the premise that a few particles are required for accurate estimation if the particles are sampled from high-probability regions of the probability distribution function. In addition, the implicit sampling technique can also mitigate the need for a resampling step to avoid filter collapse.
III Multi-agent Constrained Trajectory Games
@@
We formulate the multi-agent interactive trajectory planning problem as an $N$-agent constrained trajectory game. Let $\mathcal{N}:=\{1,2,\ldots,N\}$ denote the set of agents’ indices. We assume that the game is played over a finite time horizon of $\tau\in\mathbb{N}$ time-steps, where $\mathbb{N}$ is the set of natural numbers.
@@
In constrained trajectory games, each agent has to satisfy some constraints that couple different agents’ states and controls and, in general, depend on other agents’ states and actions. Let $g:\mathbb{R}^{n}\times\mathbb{R}^{m}\to\mathbb{R}^{c}$ with $c\in\mathbb{N}$ denote a continuously differentiable function that defines the constraints on the joint state $x_{t}$ and joint control $u_{t}$ at each time:
@@
where ${0}_{c}$ is a vector of zeros in $\mathbb{R}^{c}$. The aim of the constraint function $g$ is to capture individual agents’ strict preferences, such as constraints for each agent avoiding obstacles, or avoiding collision with other agents.
@@
$\frac{1}{2}\left\lVert x_{\tau}^{i}-\hat{x}_{\tau}^{i}\right\rVert_{Q^{i}_{\tau}}^{2}+\frac{1}{2}\sum_{t=0}^{\tau-1}\left\lVert x_{t}^{i}-\hat{x}_{t}^{i}\right\rVert_{Q^{i}}^{2}+\frac{1}{2}\sum_{t=0}^{\tau}\left\lVert u_{t}^{i}\right\rVert_{R^{i}}^{2}$ (3)
@@ III-A Local Generalized Nash Equilibrium Trajectories
Due to the interactive nature of the problem, generally, it is not possible for all agents to minimize their costs simultaneously while satisfying the constraints. Therefore, the interaction outcome is best captured by local GNE trajectories of the underlying trajectory game. With dynamics as (1) and constraints as (2), we denote the game in the compact form by
@@
For a given $\epsilon$, we define the set of local trajectories around a joint trajectory as $\mathcal{D}\left(\{{x}_{0:\tau},{u}_{0:\tau}\},\epsilon\right)=\left\{\{{x}^{\prime}_{0:\tau},{u}^{\prime}_{0:\tau}\}\bigg{|}\sum_{t=0}^{\tau}\left(\left\lVert x_{t}-{x}^{\prime}_{t}\right\rVert^{2}_{2}+\left\lVert u_{t}-{u}^{\prime}_{t}\right\rVert^{2}_{2}\right)\leq\epsilon\right\}.$ A local GNE trajectory for the game $\mathcal{G}$ is then defined as follows.
@@
$\begin{array}[]{ll}\underset{\{x_{0:\tau}^{i},u_{0:\tau}^{i}\}}{\text{minimize}}&\frac{1}{2}\left\lVert x_{\tau}^{i}-\hat{x}_{\tau}^{i}\right\rVert_{Q^{i}_{\tau}}^{2}+\frac{1}{2}\sum_{t=0}^{\tau-1}\left\lVert x_{t}^{i}-\hat{x}_{t}^{i}\right\rVert_{Q^{i}}^{2}\\ &\qquad\qquad+\frac{1}{2}\sum_{t=0}^{\tau}\left\lVert u_{t}^{i}\right\rVert_{R^{i}}^{2}\\ \mbox{subject to}&x_{0}^{i}=\hat{x}_{0}^{i},\enskip{x}_{t+1}^{i}=f^{i}({x}_{t}^{i},{u}_{t}^{i}),\enskip 0\leq t\leq\tau-1,\\ &g(\tilde{x}_{t},\tilde{u}_{t})\leq{0}_{c},\enskip 0\leq t\leq\tau,\\ \end{array}$ (4)
@@
Intuitively, at a local GNE, no agent can reduce their cost function by independently changing their trajectory to any alternative feasible trajectory within a local neighborhood of the joint equilibrium trajectory, $\{{x}_{0:\tau}^{i^{\star}},{u}_{0:\tau}^{i^{\star}}\}_{i=1}^{N}$. Note that in generalized Nash equilibria, the agents’ trajectories are coupled both through their independent objective functions and their joint constraints. It is important to highlight that solving (4) requires solving a set of $N$ coupled constrained optimal control problems, which is often computationally expensive. Therefore, finding all the local Nash equilibria of the game can be time-consuming when using traditional game theoretic methods that aim to solve coupled optimal control problems. Hence, in order to alleviate this problem, we leverage constrained potential games, which we build on to identify different modes of interaction.
IV Constrained Trajectory Potential Games
Constrained trajectory potential games are a class of games for which a GNE of the original game can be computed by optimizing a single constrained optimal control problem called Potential, as described in [13]. This reduces the computational cost of computing the GNEs of the game. In the following, we will prove that game $\mathcal{G}$ can be expressed as a constrained trajectory potential game. The following theorem shows that any local minimizer of such a potential function is a local generalized Nash equilibrium trajectory.
@@ IV Constrained Trajectory Potential Games
$\begin{array}[]{ll}\underset{\{x_{0:\tau},u_{0:\tau}\}}{\text{minimize}}&\frac{1}{2}\left\lVert x_{\tau}-\hat{x}_{\tau}\right\rVert_{Q_{\tau}}^{2}+\frac{1}{2}\sum_{t=0}^{\tau-1}\left\lVert x_{t}-\hat{x}_{t}\right\rVert_{Q}^{2}\\ &\qquad+\frac{1}{2}\sum_{t=0}^{\tau}\left\lVert u_{t}\right\rVert_{R}^{2}\\ \mbox{subject to}&x_{0}=\hat{x}_{0},\enskip x_{t+1}=f(x_{t},u_{t}),\enskip 0\leq t\leq\tau-1,\\ &g(x_{t},u_{t})\leq{0}_{c},\enskip 0\leq t\leq\tau,\end{array}$ (5)
@@
This theorem follows from Theorem 2 from [13], which states that if each agent’s cost function depends only on their own states and actions, then the underlying game is a constrained dynamic potential game. It is easy to see that agent costs according to (3) depend only on the individual agent’s own states and control inputs. Therefore, we can use the results from Theorem 2 from [13] to obtain the potential function of the game $\mathcal{G}$ to be
$\sum_{i=1}^{N}(\frac{1}{2}\left\lVert x_{\tau}^{i}-\hat{x}_{\tau}^{i}\right\rVert_{Q^{i}_{\tau}}^{2}+\frac{1}{2}\sum_{t=0}^{\tau-1}\left\lVert x_{t}^{i}-\hat{x}_{t}^{i}\right\rVert_{Q^{i}}^{2}+\frac{1}{2}\sum_{t=0}^{\tau}\left\lVert u_{t}^{i}\right\rVert_{R^{i}}^{2}).$ (6)
Furthermore, using the joint notations, we can re-write the sums in (6) as $\sum_{i=1}^{N}\left(\frac{1}{2}\left\lVert x_{\tau}^{i}-\hat{x}_{\tau}^{i}\right\rVert_{Q^{i}_{\tau}}^{2}\right)=\frac{1}{2}\left\lVert x_{\tau}-\hat{x}_{\tau}\right\rVert_{Q_{\tau}}^{2},$ $\sum_{i=1}^{N}\left(\frac{1}{2}\left\lVert x_{t}^{i}-\hat{x}_{t}^{i}\right\rVert_{Q^{i}}^{2}\right)=\frac{1}{2}\left\lVert x_{t}-\hat{x}_{t}\right\rVert_{Q}^{2},$ $\sum_{i=1}^{N}\left(\frac{1}{2}\left\lVert u_{t}^{i}\right\rVert_{R^{i}}^{2}\right)=\frac{1}{2}\left\lVert u_{t}\right\rVert_{R}^{2}$. which in turn gives us (5).
@@
Therefore, we have proven that $\mathcal{G}$ is a constrained trajectory potential game with potential (5). Hence, using Theorem 2 from [13], any local solution of (5) will also be a local GNE of the game $\mathcal{G}$. ∎

### v2

@@ Abstract
Modern robotic systems frequently engage in complex multi-agent interactions, many of which are inherently multi-modal, meaning they can lead to multiple distinct outcomes. To interact effectively, robots must recognize the possible interaction modes and adapt to the one preferred by other agents. In this work, we propose an efficient algorithm for capturing the multimodality in multi-agent interactions. We leverage a game-theoretic planner to model interaction outcomes as equilibria where each equilibrium corresponds to a distinct interaction mode. We then develop an efficient algorithm to identify all the equilibria, allowing robots to reason about multiple interaction modes. More specifically, we formulate interactive planning as Constrained Potential Trajectory Games (CPTGs) and model interaction outcomes by local Generalized Nash equilibria (GNEs) of the game. CPTGs are a class of games for which a local GNE can be found by solving a single constrained optimal control problem where a potential function is minimized. We propose to integrate the potential game approach with implicit particle filtering, a sample-efficient method for non-convex trajectory optimization. We utilize implicit particle filtering to identify the coarse estimates of multiple local minimizers of the game’s potential function. MultiNash-PF then refines these estimates with optimization solvers, obtaining different local GNEs. We show through numerical simulations that MultiNash-PF reduces computation time by up to 50% compared to a baseline. We further demonstrate the effectiveness of our algorithm in real-world human-robot interaction scenarios, where it successfully accounts for the multi-modal nature of interactions and resolves potential conflicts in real-time.
@@ I Introduction
Modern-world robotics applications are inherently multi-agent in nature. For instance, autonomous cars on highways need to account for other vehicles while planning their motion. Most of the real-world interactions are multi-modal. For example, in the two-player interaction illustrated in Fig. 1, the agents need to move towards their goals while avoiding collisions. Here, two outcomes exist: both agents yielding to their right or left. If the two agents were to act independently, they could reach different outcomes, which would result in a collision. What trajectory each agent would prefer to pick depends on the convention they might be following. For example, in some cultures, in such settings, the convention is to yield to the right to avoid collisions while in other cultures, the convention is to yield to the left. Such conventions usually arise out of recurring coordination problems [1]. Therefore, autonomous agents need to be aware of the existence of multi-modality in the interactions for efficient coordination with people who follow different conventions. Some recent works have shown that computing such multiple trajectories is critical to aligning agents’ preferences over different outcomes [2] or helping coordinate agents’ actions via recommendation [3].
Figure 1: Snapshot of an interaction between a robot and a human when they exchange positions. Using MultiNash-PF the robot can compute both modes of interaction. At $t=2$, the robot is uncertain about the preferred mode of the human since the human trajectory’s distance from both equilibria is almost the same. At $t=4$, the human trajectory is closer to the second GNE therefore, the robot chooses to follow the second GNE accordingly and avoid collision with the human.
@@
To formalize this problem, we will use dynamic game theory [4] which has shown to be powerful in developing interactive and socially complaints robots [5, 6, 7, 8]. We formulate a multi-agent interactive trajectory planning problem as a Constrained Potential Trajectory Game (CPTG). In CPTGs, the interaction outcomes are captured through generalized Nash equilibrium (GNE), which are constrained Nash equilibria. When multiple agents- such as robots and humans- interact, a GNE corresponds to a joint action such that no agent has incentives to deviate from its equilibrium actions unilaterally. Several recent works have shown great success in utilizing GNE for multi-agent motion planning in interactive domains [9, 10, 11, 12]. A key challenge in interactive domains, however, is the existence of multiple equilibria, many of which are equally valid. We argue that each equilibrium corresponds to a distinct mode of interaction. For instance, in Fig. 1, yielding to the right or left are both acceptable equilibria, representing different Nash equilibria. Consequently, we propose that in order to capture different interaction modes, robots must reason about multiple interaction equilibria and select the one that aligns with the convention adopted by other agents.
@@
Computing different local GNEs requires an efficient way to solve games. In practice, computing even a single GNE is often challenging [13] because solving for GNE amounts to solving a set of coupled constrained optimal control problems. Typically, such a solution requires solving first-order KKT conditions of the underlying problems. Therefore, solving the GNE often reduces to addressing a set of complementarity conditions, a process known to be computationally demanding, especially for large-scale applications [14]. Furthermore, computing all the different local GNEs requires initializing the game solvers with many different initial guesses of the equilibrium trajectories from prior knowledge of the specific scenario. This further increases the required number of initial guesses for the game solvers to find all local GNEs. To our best knowledge, there is no principled approach to compute multiple local Nash equilibria in constrained trajectory games other than exhaustive random initializations of the game solvers.
@@
In this paper, we tackle this challenge by introducing MultiNash-PF, a novel framework designed to efficiently compute multiple local generalized Nash equilibria (GNEs) in constrained trajectory games. This, in turn, enables robots to recognize and reason about multiple interaction modes. Our approach integrates constrained potential games [15], a class of games that allow for the computation of local GNEs by solving a single constrained optimal control problem, with constraint-aware implicit particle filtering—a sample-efficient method recently applied to nonconvex trajectory optimization [16] to achieve efficient identification of multiple equilibria.
Using CPTGs, we can compute a local GNE by solving a single constrained optimal control problem using off-the-shelf optimization solvers. Then, we use a constraint-aware implicit particle filtering method to identify multiple local minima of the potential minimization problem. This method eliminates the requirements of random initialization of the optimization solver by efficiently obtaining coarse estimates of different local GNEs. Next, we use the identified minima as initialization for an optimization solver to refine our solution. The solver then computes all the different local GNEs of the original game. Leveraging the sample efficiency of the implicit particle filtering method, MultiNash-PF can explore and identify the multiple local GNEs faster than the random initialization of optimization solvers. We further show the real-time application of MultiNashPF in a human-robot navigation setting where the robot can reason about the existence of different local GNEs and plan its motion autonomously based on the modality the human prefers.
@@
Figure 4: When using MultiNash-PF on a difficult two agent trajectory game, we first employ an implicit particle filter to obtain coarse estimates of GNEs.
@@ II Related Works
Game Theoretic Planning. One of the earliest works laying the foundation of game theoretic motion planning was [17]. Game theoretic motion planners that seek Stackelberg equilibria were proposed in [18, 7]. Iterative best response methods for computing Nash equilibria were developed in [19, 20]. Iterative linear quadratic approximations to compute Nash equilibria were introduced in [6]. ALGames was developed in [21] which is an augmented lagrangian-based solver for constrained dynamic games. A sequential linear-quadratic technique game-based technique to compute GNE was proposed in [22]. However, these methods are computationally expensive for real-time implementations as the number of agents increases. Potential games-based methods were developed later in [15, 23, 11, 12] to reduce the complexity of computing equilibria. However, the shortcomings of all these methods lie in the fact that they find only one Nash equilibrium of the underlying game and ignore the multi-modal nature of the problem that is crucial for coordination among agents. In [2], the authors recognize the existence of multiple equilibria and focus on inferring the equilibrium of other agents in the environment. However, the methodology relies on computing the different approximate local Nash by rollouts of randomly selected initial strategies, which we show in this work performs poorly compared to our method.
@@
Multiple Local Solutions to Constrained Optimal Control. In general, optimal control problems are often formulated as nonconvex constrained optimization problems, which are solved through the conventional gradient-based approaches such as inter-point methods [24] and sequential quadratic programming [25]. However, given an initialization, they can only find a single local optimal solution. However, in the present context of identifying all potential local GNEs, the gradient-based method would require numerous initializations to find all the local solutions. Departing from gradient-based methods, a Bayesian inference approach has been proposed to solve model predictive control problem [26, 27, 28]. The approach formulates the optimal control problem from a Bayesian inference perspective and employs particle filtering as its estimation method. At its core, particle filtering utilizes sequential Monte Carlo sampling to perform state estimation for highly nonlinear systems [29, 30]. The efficiency and accuracy of the estimation largely depend on the quality of the sampled particles. Recently, in [31], an efficient particle filtering method was proposed based on the principle of implicit sampling.
III Constrained Potential Trajectory Games
@@
We formulate the multi-agent interactive trajectory planning problem as an $N$-agent constrained potential trajectory game (CPTG). Let $\mathcal{N}:=\{1,2,\ldots,N\}$ denote the set of agents’ indices. We assume that the game is played over a finite time horizon of $\tau\in\mathbb{N}$ time-steps, where $\mathbb{N}$ is the set of natural numbers.
@@
In CPTGs, each agent has to satisfy some constraints that couple different agents’ states and controls. Let $g:\mathbb{R}^{n}\times\mathbb{R}^{m}\to\mathbb{R}^{c}$ with $c\in\mathbb{N}$ denote a continuously differentiable function that defines the constraints on the joint state $x_{t}$ and joint control $u_{t}$ at each time:
@@
where ${0}_{c}$ is a vector of zeros in $\mathbb{R}^{c}$. The aim of the constraint function $g$ is to capture individual agents’ strict preferences, such as each agent avoiding obstacles, or avoiding collision with other agents. It should be noted that coupling amongst agents’ decisions is captured through constraints in CPTGs.
@@
$\left\lVert x_{\tau}^{i}-\hat{x}_{\tau}^{i}\right\rVert_{Q^{i}_{\tau}}^{2}+\sum_{t=0}^{\tau-1}\left\lVert x_{t}^{i}-\hat{x}_{t}^{i}\right\rVert_{Q^{i}}^{2}+\sum_{t=0}^{\tau}\left\lVert u_{t}^{i}\right\rVert_{R^{i}}^{2}$ (3)
@@ III-A Local Generalized Nash Equilibrium Trajectories
Due to the interactive nature of the problem, generally, it is not possible for all agents to minimize their costs simultaneously while satisfying the constraints. Therefore, the interaction outcome is best captured by local GNE trajectories of the underlying trajectory game. With dynamics as (1) and constraints as (2), we denote the game by
@@
For a given $\epsilon$, we define the set of local trajectories around a joint trajectory as $\mathcal{D}\left(\{{x}_{0:\tau},{u}_{0:\tau}\},\epsilon\right)=\left\{\{{x}^{\prime}_{0:\tau},{u}^{\prime}_{0:\tau}\}\bigg{|}\sum_{t=0}^{\tau}\left(\left\lVert x_{t}-{x}^{\prime}_{t}\right\rVert^{2}_{2}+\left\lVert u_{t}-{u}^{\prime}_{t}\right\rVert^{2}_{2}\right)\leq\epsilon\right\}.$ A local GNE trajectory for $\mathcal{G}$ is then defined as follows.
@@
$\min_{\{x_{0:\tau}^{i},u_{0:\tau}^{i}\}}$ $\left\lVert x_{\tau}^{i}-\hat{x}_{\tau}^{i}\right\rVert_{Q^{i}_{\tau}}^{2}+\sum_{t=0}^{\tau-1}\left\lVert x_{t}^{i}-\hat{x}_{t}^{i}\right\rVert_{Q^{i}}^{2}+\sum_{t=0}^{\tau}\left\lVert u_{t}^{i}\right\rVert_{R^{i}}^{2}$
$\mathrm{s.t}\quad$ $x^{i}_{0}=\hat{x}^{i}_{0},\enskip x^{i}_{t+1}=f^{i}(x^{i}_{t},u^{i}_{t}),\enskip 0\leq t\leq\tau-1,$
$g(\tilde{x}_{t},\tilde{u}_{t})\leq{0}_{c},\enskip 0\leq t\leq\tau,$ (4)
@@
Intuitively, at a local GNE, no agent can reduce their cost function by independently changing their trajectory to any alternative feasible trajectory within a local neighborhood of the joint equilibrium trajectory, $\{{x}_{0:\tau}^{i^{\star}},{u}_{0:\tau}^{i^{\star}}\}_{i=1}^{N}$. Note that in CPTGs, the agents’ trajectories are coupled through their joint constraints. It is important to highlight that solving (1) requires solving a set of $N$ coupled constrained optimal control problems, which is often computationally expensive. Therefore, finding all the local GNEs of the game can be time-consuming when using traditional game theoretic methods that aim to solve coupled optimal control problems.
However, an important property of CPTGs is that a GNE of the original game can be computed by optimizing a single constrained optimal control problem called Potential, as described in [11]. This reduces the computational cost of computing the GNEs of the game. In the following, we will prove that game $\mathcal{G}$ can be expressed as a constrained trajectory potential game. The following theorem shows that any local minimizer of such a potential function is a local generalized Nash equilibrium trajectory.
@@
$\min_{\{x_{0:\tau},u_{0:\tau}\}}$ $\left\lVert x_{\tau}-\hat{x}_{\tau}\right\rVert_{Q_{\tau}}^{2}+\sum_{t=0}^{\tau-1}\left\lVert x_{t}-\hat{x}_{t}\right\rVert_{Q}^{2}+\sum_{t=0}^{\tau}\left\lVert u_{t}\right\rVert_{R}^{2}$
$\mathrm{s.t}\quad$ $x_{0}=\hat{x}_{0},\enskip x_{t+1}=f(x_{t},u_{t}),\enskip 0\leq t\leq\tau-1,$
$g(x_{t},u_{t})\leq{0}_{c},\enskip 0\leq t\leq\tau,$ (5)
@@
This theorem follows from [11, Thm. 2]. Since agent costs from (3) depend only on the individual agent’s own states and control inputs, we use the results from [11, Thm. 2] to obtain the potential function of $\mathcal{G}$ to be
$\sum_{i=1}^{N}\left(\left\lVert x_{\tau}^{i}-\hat{x}_{\tau}^{i}\right\rVert_{Q^{i}_{\tau}}^{2}+\sum_{t=0}^{\tau-1}\left\lVert x_{t}^{i}-\hat{x}_{t}^{i}\right\rVert_{Q^{i}}^{2}+\sum_{t=0}^{\tau}\left\lVert u_{t}^{i}\right\rVert_{R^{i}}^{2}\right).$ (6)
@@
Furthermore, using the joint notations, we can re-write the sums in (6) as (1). Therefore, we have proven that $\mathcal{G}$ is a CPTG with potential (1). Hence, using Theorem 2 from [11], any local solution of (1) will also be a local GNE of $\mathcal{G}$. ∎
Using Theorem 1, any local solutions of (1) will also be local GNE trajectories. We would like to acknowledge that finding local solutions of (1) will not necessarily give us all the GNEs of the original as Theorem 1 provides only a sufficient condition for computing local GNEs. Furthermore, it is known that the problem of finding all Nash equilibria is a PPAD-complete problem even in the case of static games [13]. However, for all practical purposes in multi-agent navigation, finding local solutions of (1) has proven to be enough for us.